OpenStudy (andijo76):

an online presented this question "would the recent norovirus outbreak deter you from takind a cruise?" among the 34,271 people who responded 62% answered "yes" use the sample data to construct a 95% confidence interval estimate for the proportion of the population of all the people who would respond yes to that question does the confidence interval provide a good estimate of the population proportion

OpenStudy (andijo76):

i am lost i have so far but i think its wrong a=0.1[x(0.05)=1.645 90% confidence interval is p+/- z*sqrt(p*(1-p)/n) -----> 0.62 +/- 1.645*sqrt (0.62*(1-0.62)/34,271 now this is where i am lost

OpenStudy (mathmale):

Hi, Andi, 0.62 +/- 1.645*sqrt (0.62*(1-0.62)/34,271 looks quite good; all I can say at this point is that you really need a right parenthesis just after the 34,271. 0.62 +/- 1.645*sqrt (0.62*(1-0.62)/34271) Omit the comma. Could you elaborate a bit: Why do you feel lost?

OpenStudy (andijo76):

i was following along and there is still one step or at least I thought there was

OpenStudy (andijo76):

there is according to my sample one more step i dont know if i combine the last step to get a total

OpenStudy (andijo76):

the total should be two sets of numbers that gives me the answer for _<p<_

OpenStudy (mathmale):

Good morning, Andijo: 0.62 +/- 1.645*sqrt (0.62*(1-0.62)/34271) represents a confidence interval, and should, as you yourself have pointed out, be written as _<p<_ or within parentheses, like ( a , b ). Both of these represent intervals, with a low starting value and a higher ending value. Here, that 0.62 on the left is the center of the interval, and the 1.645*sqrt (0.62*(1-0.62)/34271) part is calculated and then both subtracted from and added to the center (0.62). 1.645*sqrt (0.62*(1-0.62)/34271), called the "margin of error," is calculated as follows: 1.645*sqrt (0.62*(0.38)/34271) = 1.645*sqrt (0.24/34271) = 1.645*sqrt (0.000007) = 1.645*0.0026 = 0.004 We first subtract that from the 'center,' 0.62, and then add it to .62, obtaining (0.62-0.004, 0.62+0.004). Please complete the subtraction addition here. The result is your "95% confidence interval for the true 'proportion of the population of all the people who would respond yes to that question' (which is your original wording)" You could also write this as \[0.616 < p < 0.624.\]

OpenStudy (mathmale):

Ask any further questions about this problem that you may want to ask.

OpenStudy (andijo76):

I think i get it for now till my brain malfunctions lol thanks

OpenStudy (mathmale):

Looks complicated at first!!! But after you've done this sort of problem a couple of times, it won't look so complicated.

OpenStudy (andijo76):

ok thank you i seem to get it if i do a problem over and over a few times